you see a sudden flash light,then 6 seconds later you hear the thunder.take the speeds of sound and light as 340ms^-1and3*10^8ms^-1,respectively.how far away was the lightining strike?
1
Expert's answer
2020-04-10T13:41:16-0400

According to the equation:

t = d / v

where t - time, d - distanse, v - speed.

As light and sound passed through the same distance and difference between the times required is 6 sec, then:

(d / vsound) - (d / vlight) = 6

From here:

d (1/vsound - 1/vlight) = 6

d = 6 / (1/vsound - 1/vlight) = [6 × vsound × vlight] / [vlight - vsound]

As vlight >> vsound, the equation can be written as following:

d = [6 × vsound × vlight] / vlight = 6 × vsound = 6 s × 340 m/s = 2040 m = 2.04 km


Answer: 2.04 km